Arrange Se, K, and Ga in terms of increasing Zeff values acting on a valence electron. Assume that the valence electrons contribute nothing to the Zeff values. (Use the appropriate <, =, or > symbol to separate substances in the list.)
Answer #1

2 = 2-5 2= no ob protons s = co-ne electrons k=19= 15 252 2p6 35 3p6 45 coure electrons = 18 protona 19 Zeff = 19-18=1 Gra= 3